Operations Report – January 2nd

Crew 186 Operations Report 02JAN2018

SOL: 02

Name of person filing report: M. Grande

Non-nominal systems: Generator power system

Notes on non-nominal systems:

Ready for a long Engineering story? Following the generator failure early yesterday morning, I’ve been keeping an eye on the batteries and the battery charger. Yesterday, Sol 01, I removed the plastic covers from the posts to break out the “new” battery (with the sticker “9/17”), plugging that into the charger. As noted in Sol 01 Operations Report, I switched out the batteries in the evening, so the “old” battery (sticker “10/17”) that we had been using all the time was back in action. Though the batteries should supposedly last us 24 hours, the new 9/17 battery was unfortunately at a 50% SOC yesterday evening, after running the generator through our Martian daylight hours. This may be due to my lack of knowledge in the need to charge batteries before usage, and I would please request clarification on this requirement and possible effects of not charging before first use.

I mention this concern because today, Sol 02, has been a bright, sunny, and clear day… Perfect for spending the day taking trips through the tunnels to keep checking on an off-nominal battery charger! This morning, the battery charger showed 0% SOC after it was supposed to be charging the new 9/17 battery all night long! Too cold in the Science Dome maybe, even though I moved it inside? Also this morning, the generator indicated low battery and was running at only 9.3 V (lower than nominal 12 V). The rugged old 10/17 battery had charged all day, supposedly, but now apparently was at low charge after only running through the night.

Well, Max and I hooked up the multimeter he brought to check the voltages and amp draw on both batteries. At some point, the charger also began showing around 50% SOC on both the batteries. The charger was connected to the old 10/17 battery for several hours in the morning, and then was connected to the new 9/17 battery for several hours–in fact, all afternoon. Wrong SOC readings on the charger? The 50% generator battery SOC seems to be a small success, because the multimeter readings indicate the batteries should be full… On both batteries, now: 13.0 +/- 1.0 V and 5.5 +/- 0.6 A. Based on some battery knowledge and reading over a couple Schumacher Battery Chargers 10/2 A output, this indicated a fully charged battery. Small success?

The trusty old 9/17 battery has been plugged in this evening around 5pm when the Hab switched from solar power. The 10/17 battery was plugged into the generator at first, showing 100% SOC on the charger actually, but the generator failed to start. So I brought it back in and am leaving it plugged into the charger inside the Science Dome overnight.

We’ll continue to keep an eye on the generator system! The battery charger remains suspicious and on trial.

Generator (hours run): 12.8h

Generator turned off, charging 10/17 battery at 8:45am

Charging 9/17 battery at 12:30pm

Generator turned on at 5:50pm

Solar— SOC

@ 8:00am : 100%

@ 5:50pm : 60%

Diesel: 65%

Propane: 39%

Ethanol Free Gasoline (5 Gallon containers for ATV): 00 Gallons

Water (trailer): 10 Gallons

Water (static): 430 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 128950.1 Gallons

Toilet tank emptied: No

ATVs Used: None

Oil Added? No

ATV Fuel Used: 00 Gallons

# Hours the ATVs were used today: 00:00 hours

Notes on ATVs: None

Deimos rover used: No

Hours: 96.4

Beginning charge: 100%

Ending charge:

Currently charging: Yes

Sojourner rover used: Assigned to director only.

Hours: 5.1

Beginning charge: Unknown

Ending charge: 96%

Currently charging: No

Spirit rover used: Yes

Hours: 7.6

Beginning charge: 100%

Ending charge: 89%

Currently charging: Yes

Opportunity rover used: No

Hours: 4.3

Beginning charge:

Ending charge:

Currently charging: No

Curiosity rover used: Yes

Hours: 3.1

Beginning charge: 100%

Ending charge: 72%

Currently charging: No

HabCar used and why, where? No

General notes and comments:

Curiosity rover was charging from the end of the EVA until the Engineering EVA, when I switched the cord to Spirit. However, I forgot to check the final charge on Curiosity after the Engr EVA, my apologies. We need 1-2 more extension cords in order to have all rovers charging at one time.

I took my first real Engineering EVA this evening with Commander Max! We trekked over to check the propane, unloaded some things for the little NorCal rover from the RAM, and I checked on a couple of our electric rovers. Tried out the Exo-Suit for the first time, too! Felt great at first and definitely so easy to slip on, and I appreciated it not hitting the back of my head like the other EVA packs–poor short Mel. Halfway through (which ws only after about 10 mins) I did feel some pretty uncomfortable strain on my shoulders. I’ll try adjusting it better next time.

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support:

Maybe we can find someone with more generator and battery knowledge to check out what’s been going on. I’ve tried to record everything in my notebook and summarize them for CapCom so far. Also, I replaced the air filter, and Shannon says this should become routine at the start of each rotation from now on.

Regards,
Melanie Grande, Crew 186 Engineer

Sol Summary Report – January 2nd

Crew 186 Sol 2 Summary Report 02JAN2018

Sol 2

Summary Title: Elusive Blueberries and Baby Radishes

Author’s name: Max Fagin

Mission Status: All nominal

Sol Activity Summary: Our second day on Mars was opened with the slow-motion electrical chords of Vangelis theme from Chariots of Fire. After our morning Yoga, the second EVA team composed of Kshitij, Cesare, Mark, and Sam departed the airlock to search and sample the hematite blueberry formations at Goldstone Rd. that had eluded the first EVA team. While they were out, the remaining crew members caught up on Hab tasks, including Melanie dealing with a finicky battery charger, and Max watering our brand new Radish microgreens that Mark has spent the last few days preparing GreenHab for. The EVA team had a very successful (if exhausting) long duration EVA, stopping twice under Cesare’s guidance to sample interesting geological features. The hematite blueberries, alas, continued to evade us. Shannon: We formally request a hint.

The connection to the Hanksville radio repeater we had hoped to use today for long-range communication wasn’t ready, so communications were lost between the EVA and Hab almost as soon as they rounded the corner, but the connection was fixed by the time they had returned, and we intend to use it to keep in touch on tomorrow’s EVA. The evening finished with Sam beginning his microbe sampling swabs around the hab, and an engineering EVA to move rover supplies from the RAM, along with a brand new load of soil for GreenHab that a friendly supply shuttle dropped right on our doorstep.

Our crew has developed a new tradition which we have dubbed “The Idiot Box”. It is a box which we leave just outside the airlock all day, into which we put all the equipment which we will feel like an idiot for forgetting on our next EVA. That missing airlock fastener? That thermal probe? That EVA medical kit? Ask any one of us about any of these pieces of equipment and we will all give you the same answer: “Put it in the Idiot Box.”

Look Ahead Plan: Our first non-geology EVA will take place tomorrow, with the three goals of photographing the as yet unvisited Maxwell Montes, testing our new long-range communications system with the Hanksville repeater, and testing the direction radio beacon for navigation.

Anomalies in work: The new generator battery does not seem to be charging as expected

Weather: 0C, Winds W at 3MPH, Humidity 43%

Crew Physical Status: Healthy

EVA: Greenstone Rd take 2, 3.25 hrs, geological sampling for spectral analysis

Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, EVA Report, EVA Request, Geology Report

Support Requested: The newer of the two generator batteries that we have on hand is not charging as expected. This may simply be due to it being relatively unused (see Operations report for details), but we aren’t sure. We can continue on the one battery for now, but would appreciate advice or sources on breaking in a new battery.

EVA Report – January 2nd

Crew 186 EVA Report 02JAN2018

Author: Cesare Guariniello

Purpose of EVA: Complete EVA #1 that was cut short: spectral sampling of clays and searching for hematite spherules in Greenstone region.

Destination: Greenstone region east of Greenstone Rd

UDM27 Coordinates:
520050E
4248300N

Participants: Kshitij, Cesare, Mark, Sam

Narrative: This EVA was the first one for the three crew members who had not participated into the EVA yesterday. Since the first EVA had been cut short, the crew geologist joined this EVA to complete the tasks planned for the previous day. The planned sites were the region East of Greenstone Rd, and on the way back the region between Zubrin’s Head and Robert’s Rock Garden, both for collecting more clays and looking for hematite “blueberries”. The crew felt unusual levels of heat, due to a clear bright Sun, which also made it challenging to drive heading South, and possibly prevented the crew from identifying the access to Greenstone Road. Feeling that they might have overshot the intended destination, the crew parked the rovers and found out that they had reached White Rock Canyon, slightly South of Greenstone Road. The incredible sight and the discovery of a stream of frozen water at the bottom of the canyon prompted the EVA crew to take a few photos to document the area, before heading back to the rovers and finally reaching the intended destination, at the beginning of Greenstone Road. While the crew had no luck with the search for hematite blueberries, Cesare analyzed and collected more clay sample, with the support of Kshitij, Mark, and Sam. In the meanwhile, Kshitij also grabbed videos of the crew activity, while Sam took breathtaking photos of the location and of the crew.

After spending one hour and fifteen minutes at the first location, the crew headed back Northwest to stop at a second location, to the West of Cow Dung Road, where one more hour was spent in field activities. Cesare found a couple of promising rounded pebbles which appear to have a spectrum that at least partially resembles hematite.

Having lost communication with the habitat after crossing the first ridge, the crew observed line-of-sight rules for safety, and -among the research work- took time to think of their expedition and their goals. We were all amazed at the extreme resemblance of the location to plains on Mars, in colors, aspect, and mineralogy. Watching the landscape from the top of a ridge while wearing a space suit is an amazing experience!

When the EVA crew members came back, they were welcomed by the rest of the crew with well deserved water, before a very useful debriefing that will help in future EVAs. The crew performed very well with the 25-50-25 rule, actually spending only thirty minutes on the drive out and back, and two hours and fifteen minutes in the field!

Cesare Guariniello, PhD

Geology Report – January 2nd

Crew 186 Geology Report 02JAN2018

The plan of today’s EVA included the two locations that were not visited yesterday: the area East of Greenstone Rd and the area between Robert’s Rock Garden and Zubrin’s Head, both in the Morrison formation. During the drive outbound, the crew geologist gave a short introduction to the other EVA members, who had not participated in the first EVA.
The crew spent a long time in the field, and the geologist was able to select, analyze, and collect various samples of clays in different strata of the Morrison formation. The search for hematite spherules was only partially successful, with two potential candidates found in the second site. The geologist also recorded temperature of the samples, for later analysis of thermal inertia of the material. The problem with the portable spectrometer is currently being addressed by personnel at Purdue.

😊

Cesare Guariniello, PhD

Journalist Report – January 2nd

Crew 186 Journalist Report 2Jan2018

[Sol 2]

The weather is exceptional today. Only faint tendrils of moisture drift lazily across the boundless sky, and we were able to run the habitat for most of the day on solar power alone. The downside to all this, however, could be attested to by the members of the crew who endured the oppressively hot radiance on our second EVA.

Guided by the crew’s geologist, Cesare, the team drove far to the south in search of clay and hematite samples. At one of the sites the crew was able to ascend a nearby butte and was treated to a spectacular panorama of the endless Martian desert. They returned with an overwhelming number of photos and much exhaustion. Our Green hab scientist, Mark Gee, had an especially tiring day, as he spent much of the morning preparing an experiment growing microgreens in conditions of minimal input and testing the effects of the habitat biome on plant growth.

Following a brief rest, the crew returned to preparing experiments and performing maintenance duties on the habitat. These duties included an engineering EVA to move equipment and check the battery status the rovers.

Justin Mansell, MDRS Crew 186 Journalist

Journalist Report – January 1st

[Sol 1] [New Year, New Planet]

The first sunrise of 2018 broke the horizon at approximately 7:40 am this morning and lit the sky with a fiery glow not unlike the sea of ochre shades below. The crew roused with little hesitation and started the day with a yoga session led by our executive officer, Kshitij Mall. By 9 am the team had donned our EVA suits and capitalized on the morning light to conduct photo shoot. It was also our last chance to enjoy the outdoors while still being able to take our helmets off.

With the crew portraits obtained, the crew returned inside the habitat to indulge in a delicious New Year breakfast of fruit crêpes, cooked for the team by yours truly. At noon, the airlocks were shut from both sides and the simulation began.

But our crew isn’t one to vegetate inside the habitat (as cozy as it can be). Almost as soon as the simulation had begun, our ardent commander, Max Fagin, had 4 of us suiting up for our first EVA. The purpose was to analyze rock samples a short distance south of the hab and the EVA also served as an important test of our spectrometer and radio navigation equipment. Upon our return, the crew engineer, Melanie Grande, used some surplus time to replace the window on the outer door of the southern airlock.

Though the EVA was short, it has given the team much to think about. We have since been discussing various logistical items and how to iron out minor bugs in our EVA equipment.

Justin Mansell, MDRS Crew 186 Journalist

P.S. Daily photos attached. Picture of the day: 01Jan2018 Happy New Year.jpg

Operations Report – January 1st

Crew 186 Operations Report 01JAN2018

SOL: 01

Name of person filing report: M. Grande

Non-nominal systems: None

Notes on non-nominal systems: Generator system limping along with a now-nominal routine.

Generator (hours run): 25h 30min

Solar— SOC

@ 9:50am : 61%

@ 5:10pm : 100%

Diesel: 70%

Propane: 40%

Ethanol Free Gasoline (5 Gallon containers for ATV): 0 Gallons

Water (trailer): 10 Gallons

Water (static): Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 128894.7 Gallons

Toilet tank emptied: Yes

ATVs Used: 350.1, 350.2

Oil Added? Yes

ATV Fuel Used: 0.1 Gallons

# Hours the ATVs were used today: 01:00 hours

Notes on ATVs: ATVs were nominal. They all need an oil change, especially the Yamaha 300, because the oil during my check was quite dirty and black. All got some oil added to them except the 300.

Today I went on an Engineering round to do a quick inspection of the ATVs: checked the oil and fuel levels, attempted to check tire pressure, and did a visual tire inspection. All results were nominal and nothing unexpected; however, I considered the quick inspection my due diligence as an Engineer on their first day! It was kind of entertaining for me anyway. Oil levels were about half-way so I added some from our stores (plenty of bottles of oil in the EVA storage room). I wanted to check the tire pressures, but unfortunately the little pen gauge is inoperational and the air compressor has a gauge without enough detail for me to get an accurate reading. All the tires seem in great condition though, and I looked for possible damage to the tire walls from going over a bad bump or something out here, too. Lookin’ good!

Deimos rover used: No

Hours: 96.4

Beginning charge: 100%

Ending charge:

Currently charging: Yes

Sojourner rover used: Assigned to director only.

Hours: 4.9

Beginning charge:

Ending charge:

Currently charging: Yes

Spirit rover used: Yes

Hours: 6.8

Beginning charge: 100%

Ending charge:

Currently charging: Yes

Opportunity rover used: No

Hours: 4.3

Beginning charge: 100%

Ending charge:

Currently charging: No

Curiosity rover used: No

Hours: 4.6

Beginning charge: 100%

Ending charge:

Currently charging: Yes

HabCar used and why, where? No

General notes and comments:

Sol 0 and as the Crew Engineer, I went around the hab for my checks while also looking things over to create a To Do list for my couple weeks here. Toilet got emptied (such fun), ATVs got a quick
inspection, and the front window replacement and back-up generator relocation was fit into the end of our EVA. I got confirmation from Shannon that we could try hooking up power to the RAM using the little back-up generator, so two crew were very helpful in wheeling it through the sand over there. We’ll just wait on the 30A cable to be delivered in a resupply mission soon, along with tools for the RAM! Yay! Finally, I also gave a little engineering brief to the crew–what to expect while they’re here with things like doors that don’t work properly and staying away from the power system and helping me record rover runtime hours.

A notable thing did happen today in that the generator failed sometime this morning. I’m not entirely sure what time, because the clock is not correct on it. Unfortunately, this means that I lost the runtime hours, too, so I can’t be sure how long it has been on. Fortunately, the Solar SOC showed 61% at that time, which is above our limit for concern. We are guessing that the battery died and that’s all. Because yesterday was cloudy and today might have been too, I left the generator on all day today but replaced the old battery with a new one. I switched the batteries again this evening, so now the new one is charging.

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support: ATVs need an oil change soon, please. Also for my reports, I requested a switch to my gmail because I believe this email server is adding additional spaces to my reports for some reason. Sorry about that!

Happy 1st of the year!

Melanie Grande, Crew 186 Engineer

Greenhab Report – January 1st

GreenHab Report

Mark Gee

01Jan2018

Environmental control: Heating with shade cloth on

Working Hour: 05:50 PM
Inside temp at working hour: 17 C
Outside temp during working hours: 6 C
Inside temperature high: 28 C
Inside temperature low: 15 C
Inside humidity: 42 %RH

Inside humidity high: 48 %RH
Inside humidity low: 25 %RH

Hours of supplemental light:

17:00 to 24:00

Changes to crops: Below is a formatted inventory of all of the crops, quantity, growth stage, actions taken, and notes. The table is also attached incase it is not visible here.

Name

Quantity

Growth Stage

Action

Notes

Dill Weed

2ft row, ~40 plants

Seedling, 3 true leaves

Rosemary

2ft row, ~20 plants

Seedling, 1 true leaf

Parsley

2ft row, ~50 plants

Seedling, cotyledon

Cilantro

2ft row, ~40 plants

Seedling, 1 true leaf

Oregano

2ft row, ~100 plants

Seedling, cotyledon

Sage

2ft row, ~10 plants

Seedling, 1 true leaf

Basil

2ft row, ~40 plants

Seedling, 1 true leaf

Thyme

2ft row, ~30 plants

Seedling, 1 true leaf

Chives

2ft row, ~20 plants

Seedling, 1 leaf

Spinach, Bloomsdale

2 seedling trays, 10 plants

Seedling, cotyledons

Need transplanting

Kale, Blue Curled Scotch

1 seedling tray, 2 pots, ~50 plants

Seedling, cotyledons

Need thinning

Cabbage, Golden Acre

1 seedling tray, ~20 plants

Seedlings, cotyledons

Need thinning

Moringa Olifera

14 plots

No plants

These are trees. Should they be grown in the small Greenhab?

Paperwhites

3 pots, seven plants

Various, sprouted to flowering

Moved flowering plants to the habitat to improve morale

Beans, Pole

27 plants

3ft vines, producing flowers and pods

Harvest in 1 week

Melon

8 plants

2ft vines, no flowers

Peppers

9 pots, 23 plants

8 inches, vegetative

Tomatoes

24 pots, 54 plants

6in-48in tall, some flowering

Need transplanting

Daily water usage for crops: 8 gallons

Time(s) of watering for crops: 05:50

Research observations: None

Changes to research plants: None.

Aquaponics: Not Functional

Narrative:

The crew entered simulation at noon. We will have to adjust to the isolation from society, cramped quarters, and space supplies. I’m hoping to make this transition easier with the produce from the Green Hab. Several of the crops are nearing harvest and I hope to cook a meal with fresh food later in the rotation to boost crew morale. For dinner, I brought some of the blooming paperwhites to the table.

I cleaned the greenhouse and packed the aquaponics equipment to make room for the microgreen experiment and have been sterilizing the soil, growing surfaces, and water.

Support/supplies needed: 3 agar plates to check seed sterilization procedures

Sol Summary – January 1st

Crew 186 Sol 1 Summary Report 01012018

Sol 1

Summary Title: New Year New Planet

Author’s name: Max Fagin

Mission Status: All nominal

Sol Activity Summary: After staying up until midnight last night to bid farewell to 2017, we awoke at 7:30AM to the music of the Launch Theme from Apollo 13. Kshitij led us in a Yoga routine that will become a daily occurrence for the rest of the mission. We had planned to finish our official crew portraits the previous evening, but sunset intervened, so we quickly donned the EVA suits this morning and stepped outside to capture our official crew portraits. On display were the flags of Purdue (our Alma Mater), our home countries, and the Tricolor-RGB of our new home planet.

Breakfast was crepes smothered with Nutella and maple syrup, prepared by Justin and Mark. Over breakfast, we established the rules of our simulation, and what was expected of each of us to maintain its fidelity. With 15 minutes left until noon (when we would officially enter simulation), we all took the time to pay one last visit to the outside. At noon, we officially entered our simulation and began preparation for our first EVA: A geological sampling mission to the Greenstone Rd. area. Entering simulation late meant we had to return to the hab after only collecting one of the targeted 3 samples, but we will return tomorrow to complete it. Lunch was leftovers from last night’s new year’s feast (it was good enough to enjoy twice), followed by a dinner and witnessing a marvelous supermoon rise. Lunar astronomy on the supermoon is planned for tonight, as the weather looks clear.

We are not the only ones who are feeling the Martian vibe. Even the habitat seems to be feeling it. For example, the generator’s 60Hz AC power seems to be running at ~61.6 Hz for some unknown reason. This is a harmless anomaly, but it means any appliance that gets its clock cycle from the AC power line will run 2.8% fast, accumulating an extra ~38 minutes per day. In other words, due to a technical error, our clocks are actually keeping within 0.2% of Martian time, and only the battery power analog clock on the wall is keeping Earth time… We have no intention of fixing this.

Look Ahead Plan: Tomorrow’s EVA will target the two sample sites we were not able to visit today. Will also experiment with the use of the Hanksville repeater (aka, communications satellite) and our ham radio to avoid comm dropouts between EVA team and MDRS. Thank you comm for the warning this morning about the solar storm! We will deploy our own telescope tonight and plan for some solar astronomy tomorrow morning.

Anomalies in work: EVA Airlock window was installed at the end of the EVA, but is currently secured with zip ties. Looking for a better way.

Weather: Temperature: 6C, Mostly clear, Winds calm, Humidity 39%, Barometer 30.27 inHg

Crew Physical Status: Healthy

EVA: Greenstone Rd, 1.0 hrs, geological sampling for spectral analysis

Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, EVA Report, Science Report (Geology)

Support Requested:

1) We installed a new EVA airlock window at the end of today’s EVA. Are there any old photos of the EVA airlock window we can use for reference to see how it is supposed to be secured? It is currently zip tied in place, which is adequate, but not pretty.

2) Can Mission Support please change the email of Crew Engineer Melanie Grande to melgrande2@gmail.com we think this will solve some formatting anomalies with the operation report.

EVA Report – January 1st

Author: Max Fagin

Purpose of EVA: Spectral sampling of sandstone clay, mudstone clay and Hematite in Goldstone region.

Destination: Between URC South and Kissing Camel Ridge, 200-300m west of Cow Dung Rd.

UDM27 Coordinates:

518100E

424850N

Participants: Max, Melanie, Cesare, Justin

Narrative: We have heard reports of hematite blueberry formations near the Goldstone Road area, and wanted to take samples with Cesare’s handheld spectrometer. Three sample sites were planned along the way to sample sandstone clays and mudstone clays. A late start meant we had to turn back after only one sample site, and limit the EVA to 1.0 hr instead of 3.5 hrs, but we made productive use of the time. While Cesare was taking spectra of rock and clay samples, Justin broke out his directional radio beacon to experiment with locating the hab. Our sample site was in a radio shadow of the the hab, so Melanie climbed a nearby hill to relay instructions between us and the Hab. This problem won’t go away on Mars (where the radio horizon is 50% closer than Earth’s), and we’ve brought equipment to mitigate it. Our crew includes three licensed ham radio operators, and tomorrow, we will use the Hanksville radio repeater (aka, “communication satellite”).

After returning, we checked two items off our engineering EVA checklist: Moving the small generator to the RAM in preparation for powering it up tomorrow, and installing a new window in the Hab airlock.

Our goal for EVA’s on this mission is to hold to a 25-50-25 rule. I.e. No more than 25% of the time from egress to ingress be spent traveling to and from to the destination, and at least 50% of the time spent on site. Today, we only spent 25 minutes on site (plus 7 minutes at the hab working on engineering tasks), which didn’t meet our goal. But we will track our performance on each future EVA to hold ourselves to improvement. Field science means field science. Not
driving-to-the-field science.